# Continual Memorization of Factoids in Large Language Models

### v1

@@ 7 Conclusion
In this paper, we formalize finetuning a language model with factual knowledge in the continual memorization framework. In contrast to continual learning, which focuses on general capability, we focus on the specific challenges inherent to finetuning on long-tail factoids. Through careful experiments, we establish that the long-tail factoid data is the largest culprit in making a model forget memorized factoids from previous stages of finetuning. We then evaluate experience replay methods that are often used in continual learning and find that they do not satisfactorily revive forgotten factoids. To address the issue of forgetting, we propose a surprising yet effective strategy REMIX. By mixing random word sequences or generic pretraining data into different stages of training, REMIX outperforms replay-based methods in our experiments despite not using any factoids from the original set in its mixing process. Finally, we analyze REMIX using Logit Lens and ablation studies to find that it teaches the model to reduce inter-fact interference by changing where it stores facts. REMIX opens up many new directions for future research. For example, future work may explore REMIX and similar approaches to ensure that safety-tuning is not easily undone by further finetuning. Its efficacy also poses interesting questions about the dynamics of memorization in language models, which we are excited to see investigated in future work.
8 Acknowledgment
We thank Alexander Wettig, Howard Yen, Tianyu Gao, and Mengzhou Xia for their valuable feedback on the manuscript. This research is also funded by the National Science Foundation (IIS-2211779) and a Sloan Research Fellowship.
@@ B.1 Main Results with Standard Deviation
B.2 Replay Results
@@
B.3 Forgetting in Familiar Factoid Instances
@@
B.4 Training Details
@@
B.5 Dataset Examples and Prompts
@@
B.5.1 Factoid dataset ($D_{A}$) Examples
@@
B.5.2 Mixing data ($D_{M}$) Examples

### v2

@@ 7 Conclusion
In this paper, we formalize finetuning a language model with factual knowledge in the continual memorization framework. In contrast to continual learning, which focuses on general capability, we focus on the specific challenges inherent to finetuning to memorize knowledge. Through careful experiments, we establish that finetuning on factoid data causes the most severe forgetting on the memorized factoids from previous stages of finetuning. We then evaluate experience replay methods that are often used in continual learning and find that they do not satisfactorily revive forgotten factoids. To address the issue of forgetting, we propose a surprisingly effective strategy REMIX. By mixing random word sequences or generic pretraining data into different stages of training, REMIX outperforms replay-based methods and other baselines in our experiments despite not using any factoids from the original set in its mixing process. Finally, we analyze REMIX using Logit Lens and ablation studies to find that it teaches the model tochange where it stores facts—moving it to earlier layers or diversifying the knowledge storage location. Studying the continual memorization problem opens up many new directions for future research. For example, future work may explore REMIX and similar approaches to ensure that safety-tuning is not easily undone by further finetuning. Its efficacy poses interesting questions about the dynamics of memorization in language models, which we are excited to see investigated in future work.
Acknowledgments
We thank Alexander Wettig, Howard Yen, Tianyu Gao, Mengzhou Xia, and Yihe Dong for their valuable feedback on the manuscript. This research is supported by the National Science Foundation (IIS-2211779), a Sloan Research Fellowship, and Cisco Research.
@@ B.1 Main Results with Standard Deviation
B.2 Mistral Results
We report the complete results of Mistral-7B-v0.3 in Table 6.
Factoid Non-Factoid
LAMA EQ WQ Avg GSM8K MATH EC APPS UC Avg
Key-Value Recall
No Mixing   0.1 15.4 29.6 15.0   4.8   1.5 12.7 13.1 51.9 16.8
Random / K-Pile 47.5 44.1 39.0 43.5 60.1 39.1 52.9 54.8 81.0 57.0
PopQA
No Mixing 66.9 92.3 89.6 82.9 96.9 96.8 96.9 96.9 96.7 96.8
Random / K-Pile 90.5 92.3 89.0 90.6 91.7 91.6 91.8 91.9 91.3 91.7
TriviaQA
No Mixing 71.6 86.4 91.5 83.2   4.8 99.0 95.9 79.9 97.0 75.3
Random / K-Pile 77.0 81.5 83.1 80.5   1.6 91.1 95.3 97.7 90.7 75.3
B.3 Replay Results
@@
B.4 Forgetting in Familiar Factoid Instances
@@
B.5 Probing Results
B.6 Ablations
Ablating mixing data length.
Figure 10: y-axis is the accuracy (%) on Key-Value Recall of varying sequence length with the mixing datasets. Top: Random Word Sequence (mixed at stage 1). Bottom: Knowledge Pile (mixed at stage 2).
Effect of mixing ratio.
Figure 11: Mixing ratio ablation. x-axis indicates the ratio of the mixing data against the training data. y-axis indicates the accuracy (%) on Key-Value Recall. The two left-most plots are both stage 2 mixing (S2) and the right-most two are both stage 1 mixing (S1).
Impact on downstream performance.
Intuitively, adding random word sequences might risk disrupting capabilities in other domains. We evaluate the model’s performance on MMLU (Hendrycks et al., 2021a) shown in Table 10. We observe REMIX maintains better performance across the board compared to no mixing.
KVR PopQA TriviaQA
No Mixing 18.5 19.0 17.5
REMIX (R / K-Pile) 24.1 27.0 21.5
Table 10: Accuracy on MMLU. We compare the No Mixing baseline to REMIX, which mixes with Random Word Sequence (R) at stage 1 and with Knowledge Pile (K-Pile) at stage 2.
Effect of different mixing data.
Figure 12: Comparison between Knowledge Pile and other generic mixing data sources: ArXiv Pile and FineWeb on KVR. y-axis indicates the accuracy (%) on KVR.
B.7 Training Details
@@
B.8 Dataset Examples and Prompts
@@
B.8.1 Factoid dataset ($D_{A}$) Examples
@@
B.8.2 Mixing data ($D_{M}$) Examples